# On Efficient Variants of Segment Anything Model: A Survey

### v3

@@ 3.1.5 Code Refactorization
The Segment Anything Fast model (SAMfast), developed by the PyTorch team [154], is a rewritten version of SAM that leverages pure, native PyTorch optimizations. SAMfast is reported to be 8x faster than the original implementation while maintaining nearly the same accuracy. This improvement is the result of a systematic process of identifying bottlenecks and applying targeted optimizations. Initially, the team identified long function calls that caused synchronous blocking, leading them to rewrite the corresponding code. Another significant bottleneck was the time-consuming matrix multiplication, which was mitigated by using bfloat16 precision. Following these adjustments, the team utilized torch.compile to fuse smaller operations and adopted PyTorch’s scaled_dot_product_attention (SDPA) to accelerate attention computations on the GPU. Additionally, by integrating the new kernel built with Triton, memory usage on the GPU was further reduced. When SAM uses the batch prediction method, input tensors of varying sizes are unified into NestedTensors, which significantly improve throughput. Despite these optimizations, matrix multiplications remained a key bottleneck. To address this, the team implemented int8 quantization and approximated matrix multiplication using semi-structured sparsity. For more details on the step-by-step optimization process, it is recommended to go through the official blog for more details.
@@ 3.1.6 Discussion
Combining different architectural components within SAM variants might further overcome the above deficiencies by complimentarily taking advantages from each architecture, leading to better balance between efficiency and accuracy. Intuitively, hybrid designs that integrate CNN-based structures with Transformer-based modules or Transformer-alternative models are feasible solutions. There have been attempts in recent SAM variants, such as Lite-SAM [124], which employs a CNN-Transformer hybrid encoder LiteViT, and RWKV-SAM [136], which integrates CNNs with the RWKV model. These variants with hybrid architectures have demonstrated significant improvements in efficiency while maintaining comparable segmentation performance to vanilla SAM. Despite their promise, the mismatch between heterogeneous layers, which may lead to low quality in feature representations and high difficulty in model optimization, is a non-negligible challenge for building up effective hybrid models, and as a result, more advanced designs about module combination are needed in the future research.
In addition to backbone selection, additional compression techniques such as quantization and pruning can further reduce model size and memory usage during inference. However, these methods may not always deliver the expected runtime speedup and can sometimes lead to a significant degradation in accuracy and generalization. The training procedure itself also has a profound impact on performance. Training from scratch can yield comprehensive, high-quality models but requires substantial training resources. Knowledge distillation offers a more efficient alternative using less data. Yet, distilled models may fail to fully inherit the teacher’s capabilities, resulting in performance inferior to models trained from scratch. Moreover, combinations of these methods may yield models with varied accuracy–efficiency profiles. For example, pruning combined with post-quantization can compress a model to an extremely lightweight variant by reducing both parameter size and model complexity; subsequent fine-tuning (either through full training or distillation) can then help recover lost accuracy, thereby achieving a more balanced design.
@@ 3.2 Accelerating SegEvery Tasks
As described in Section 3.1, the primary efficiency bottleneck for the SegAny task lies in the heavy image encoder. Any SAM variant with a more lightweight architecture will inherently be able to segment everything faster than the original SAM. However, as analyzed by Zhang et al. [155], the main challenge of the SegEvery task stems from the dense grid sampling strategy. This strategy first predicts numerous masks based on a point grid and then selects valid masks, which is computationally expensive. Consequently, designing a more efficient sampling strategy to reduce the number of predicted masks has become the core approach to accelerating the SegEvery task. Another potential solution is to convert the SegEvery task into another well-established task, such as all-instance segmentation, as done in FastSAM [41]. In this part, we will review works that have specifically proposed optimized sampling strategies to accelerate the SegEvery task.
@@
Building on the structure of SAM, Zhang et al. [155] introduced an object-aware prompt sampling strategy to enhance the efficiency of the SegEvery task. This project, named MobileSAMv2, operates independently of their previous work [43] focused on accelerating the SegAny task. In MobileSAMv2, the researchers employ the YOLOv8 model, trained on a small subset of SA-1B, for object discovery. This model generates a large number of bounding boxes corresponding to latent objects. Highly overlapping boxes are filtered using Non-Maximum Suppression (NMS), and the remaining boxes are used as box prompts. By using these filtered boxes as prompts, MobileSAMv2 eliminates the need to filter predicted masks—a much more time-consuming process. With the maximum number of prompts set to 320, the new strategy is reported to be 16 times faster than the traditional 32*32 grid sampling strategy. Additionally, MobileSAMv2 can be integrated with MobileSAM to create a unified model that achieves high efficiency in both SegAny and SegEvery tasks.
@@
Figure 14: The pipeline of hierarchical sampling strategy for TinySAM [44]. The hierarchical strategy contains two rounds of sampling: one round for sparse sampling and the other round for dense sampling (including removing points from high confident regions). The final prediction comes from the fusion of results in two rounds.
@@
Meanwhile, Shu et al. [44] also observed that using a dense points grid (e.g., 32*32, 64*64) often generates a large number of redundant masks, which are later filtered out during post-processing, an operation that incurs significant time costs. In reality, only a few points within the grid are necessary to produce confident masks. To address this inefficiency, they proposed a hierarchical strategy for efficient sampling, which progressively selects optimal points for mask generation. This strategy involves two rounds of prompt generation. In the first round, a sparse grid is used, incorporating only a fraction (approximately 1/4) of the default points along each side. Masks are generated based on these points, and after filtering, only high-confidence masks are retained as final predictions. In the second round, a denser grid is applied, following the default configuration. However, points located in regions already covered by confident masks are excluded, significantly reducing the number of points. The predicted results from both rounds are then fused to produce the final output. The pipeline for this hierarchical strategy is illustrated in Fig. 14. By employing this two-round approach, the sampling process becomes both more time-efficient and fine-grained, leading to a notable acceleration in the SegEvery task with minimal performance degradation.
@@
Figure 15: The architecture of Lite-SAM [124]. It takes the Lite-ViT as an image encoder and proposes an automated prompt proposal network (AutoPPN) for efficient sampling, while the prompt encoder and the mask decoder are copied from the original SAM.
@@
Different from all aforementioned works, Fu et al. [124] propose an end-to-end training pipeline specifically designed for the SegEvery task, aiming to develop a SAM variant capable of segmenting everything more efficiently. Their model, named Lite-SAM, retains the overall architecture of the original SAM but replaces the heavy image encoder with a more lightweight solution. The architecture overview of Lite-SAM is shown in Fig. 15. Lite-SAM employs a CNN-Transformer hybrid structure called Lite-ViT, which consists of four stages with 2, 2, 6, and 2 Lite-ViT blocks, respectively. The key innovation in Lite-ViT is the Multi-Scale Pooling Module (MSPM), which serves as an alternative to the traditional attention mechanism. Adapted from the PoolFormer block [123], MSPM utilizes channel-wise LayerNorm and extends the pooling operation to multiple scales. As discussed earlier, another major bottleneck in SAM lies in the time-consuming grid sampling strategy. To address this, Lite-SAM introduces an Automated Prompt Proposal Network (AutoPPN) to improve sampling efficiency. AutoPPN takes the feature maps generated by the encoder as input and directly predicts point and box prompts. To ensure high-quality prompts, Lite-SAM uses a more powerful MSPM-based network instead of CNNs, and incorporates distance transform to estimate the confidence of point prompts. While Lite-SAM is primarily designed to accelerate the SegEvery task, it also demonstrates improved efficiency in the SegAny task due to its lightweight image encoder.
@@
Similar to Lite-SAM’s idea of automatic prompt generation, the recent variant AoP-SAM proposed by Chen et al. [156], employs an efficient Prompt Predictor module to identify the optimal regions for prompt candidates, followed by an Adaptive Sampling and Filtering (ASF) module to generate prompts in a coarse-to-fine manner. The architecture of AoP-SAM is illustrated in Fig. 16. The Prompt Predictor consists of two encoder, processing original images and embedding from SAM’s image encoder respectively, and one decoder, incorporating the spatial and context information from encoders to generate the prompt confidence map. The function of ASF module is to extract essential prompts from prompt confidence map, which firstly samples the local maxima as coarse prompts, and secondly filters by the IoU scores of previously generated referenced masks and elimination scores of prompts, ensuring only crucial prompts will be retained for later mask generation. The elimination score is calculated by referenced mask and image embedding, indicating the possibility of being redundant. The experiments coudcted on three image segmentation datasets, have demonstrated AoP-SAM’s highly efficiency for SegEvery task, surpassing both SAM’s dense grid sampling strategy and MobileSAMv2’s [155] object-aware sampling strategy.
@@
Figure 16: The architecture of AoP-SAM [156]. It adopts a Prompt Predictor to predict prompt confidence map and an Adaptive Sampling and Filtering (ASF) module to generate final prompts.
@@ 3.3 Future Research Directions
Exploring Advanced Architectures. While current SAM variants have demonstrated efficiency gains through the adoption of efficient architectures and model compression techniques, there is substantial potential for further improvement. The exploration of Transformer-alternative models, such as Mamba [133], RetNet [193], KAN [194], and TTT [195], presents an exciting opportunity to design more lightweight and efficient structures. With the success of incorporating RWKV into SAM [136], leveraging Transformer-alternative models as backbone has shown great potential to build up efficient SAM variants, which may offer advantages in computational efficiency without sacrificing segmentation accuracy. In addition to alternative models, refining attention modules’ efficiency remains critical for further development of accelerating SAM, inspired by the preliminary exploration of FastSAM3D [128]. Methods such as linear attention, low-rank factorization, or hybrid architectures combining convolutional and attention-based designs, which have been widely applied to Transformer-based models [196, 163, 197], should be further investigated for efficient SAM. Addressing bottlenecks in both computation and memory usage will enhance SAM’s deployment across diverse hardware environments.
@@
Exploiting Sparsity and Acceleration Techniques. The inherent sparsity observed in deep neural networks, where only a subset of parameters significantly contributes to model output, offers a promising avenue for improving SAM’s efficiency. The discovery in very recent work, SuperSAM [153], showing that the subnetwork with 7 layers in image encoder pruned (12 layers in total) can still maintain over 60% performance compared to vanilla SAM, further provides proofs to SAM’s sparsity, which poses a promising research direction for further uncovering the fine-grained sparsity in SAM’s architecture. Techniques like pruning, quantization, and structured sparsity [154] could further reduce SAM’s computational demands. While several initial efforts in sparsification, such as those leveraged in SlimSAM [48] and PQ-SAM[152], have shown success, future research could dive deeper into understanding the distribution and dynamics of sparsity in SAM. This includes investigating the optimal layers or components of SAM that can be pruned or sparsified without impacting performance [153]. Additionally, techniques such as sparse attention mechanisms, dynamic pruning during inference, and low-precision training, which have been deployed on Transformer-based architectures [198, 199, 200], can be further explored to balance accuracy and efficiency, especially for large-scale deployments. By combining these with advanced knowledge distillation techniques, more compact, efficient variants of SAM could be realized.
@@
Hardware-Specific Optimizations. Optimizing SAM for specific hardware platforms, including GPUs, TPUs, specialized AI accelerators (e.g., NVIDIA’s TensorRT or Google’s Edge TPU), and edge devices, can unlock significant gains in performance and efficiency. Equipping SAM with hardware-aware model optimization techniques, for instance, operator fusion, quantization-aware training, and custom CUDA kernels, some of which have been applied in SAMfast [154] by PyTorch extensions, requires further exploration in the future to maximize throughput and reduce latency when deploying SAM on modern hardware platforms. In the context of edge devices, which face extreme constraints in storage, computational power, and energy supply, these optimizations are crucial for real-time applications, such as segmentation on UAVs or IoT devices. Future research could explore hierarchical cloud-edge architectures to offload computationally expensive tasks to the cloud, while running lightweight models locally on edge devices. This idea has already been investigated in LLMs for robot control [201]. Additionally, leveraging specialized AI hardware like Field-Programmable Gate Arrays (FPGAs) [202] or using techniques such as hardware-aware neural architecture search (NAS) and mixed-precision quantization [203] might further optimize SAM for low-latency, resource-constrained environments, ensuring that the model operates effectively across diverse hardware platforms. Among them, NAS is now successfully applied to SAM, building up the first supernetwork of SAM [153].
@@
Computation Cost Attacks and Defenses. Substantial progress has been made by the research community to improve the efficiency of deep learning models, particularly for deployment on resource-constrained edge devices. However, recent studies have exposed vulnerabilities in some models when subjected to computation cost attacks, which deliberately increase resource consumption and processing time far beyond normal operational levels [204, 205, 206, 207, 208]. Despite these findings, the efficiency and robustness of Segment Anything Model (SAM) variants under computation cost attacks remain largely unexplored, especially when techniques like dynamic models are applied. This presents a valuable opportunity for future research. One direction could involve identifying specific vulnerabilities in SAM variants when exposed to such attacks. Additionally, the insights gained from these vulnerabilities could inform the development of defense strategies that ensure stable efficiency, particularly in resource-limited environments such as edge and mobile devices.
@@
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [34] for video segmentation and MM-SAM [209] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation [210] and cross-modal feature sharing [211] could significantly reduce computational costs. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention [212], efficient memory mechanisms (which are now effective solutions for accelerating SAM 2 [213, 214]) for temporal data, and early-fusion techniques [215] to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.
@@
Efficient Variants of SAM 2. Segment Anything Model 2, successor of Segment Anything Model, which not only achieves higher accuracy and efficiency on image segmentation task but also extends its powerful capability into video segmentation tasks, has been increasingly gaining attention from industry and research community. However, similar to SAM’s dilemma, SAM 2 also suffers the inefficiency issue due to the large hierarchical image encoder, hindering its application in various resource-constrained, real-world scenarios. Moreover, another main bottleneck of SAM 2 lies in the newly introduced memory mechanism, more specifically, the memory attention module which attends feature map to memories, as memory features holds the same spatial size to image feature, resulting in long memory token sequence and heavy computational cost in cross-attention. Several works of efficient SAM 2 [216, 213, 214] has emerged focusing on solving these two bottlenecks. Surgical SAM 2 [216] has proposed the efficient frame pruning strategy which only retains the most informative frames in memory bank to achieve reduction in memory usage and acceleration in inference, providing a promising idea for efficient segmentation in videos. The subsequent variant, named EfficientTAM [213], has built up an efficient memory module, which reconstructs memories by firstly adopting average pooling in local window and secondly restoring the spatial size through repeating, leading to less computation in memory attention. Furthermore, EfficientTAM replaces the image encoder with SAMI [109] pre-trained vanilla ViT, achieving higher efficiency and maintaining competing performance. Recently, a mobile-friendly SAM 2 variant, named EdgeTAM [214], further speeds up the memory attention by leveraging a global Perciever [217] and a novel 2D Spatial Perceiver to effectively compress memories. Moreover, EdgeTAM opts for RepViT as backbone and retrains through a two-stage distillation pipeline with SAM 2 as teacher model, pursuing for efficiency and accuracy improvements. Quantization techniques have also been preliminarily employed in SAM 2, for instance, the newly proposed variant, Q-SAM2 [218], which introduces calibration in SAM 2’s image encoder and subsequently implements quantization-aware training pipeline, effectively compressing model into low-bit configurations. Compared to SAM, there are much less works on efficient variants for SAM 2 up till now, posing it a challenging task. Techniques and strategies proved to be effective in accelerating SAM, have not been fully investigated on SAM 2, which might be the potential direction for future research.

### v4

@@ 3.1.5 Code Refactorization
The Segment Anything Fast model (SAMfast), developed by the PyTorch team [SAMfast], is a rewritten version of SAM that leverages pure, native PyTorch optimizations. SAMfast is reported to be 8x faster than the original implementation while maintaining nearly the same accuracy. This improvement is the result of a systematic process of identifying bottlenecks and applying targeted optimizations. Initially, the team identified long function calls that caused synchronous blocking, leading them to rewrite the corresponding code. Another significant bottleneck was the time-consuming matrix multiplication, which was mitigated by using bfloat16 precision. Following these adjustments, the team utilized torch.compile to fuse smaller operations and adopted PyTorch’s scaled_dot_product_attention (SDPA) to accelerate attention computations on the GPU. Additionally, by integrating the new kernel built with Triton, memory usage on the GPU was further reduced. When SAM uses the batch prediction method, input tensors of varying sizes are unified into NestedTensors, which significantly improve throughput. Despite these optimizations, matrix multiplications remained a key bottleneck. To address this, the team implemented int8 quantization and approximated matrix multiplication using semi-structured sparsity. For more details on the step-by-step optimization process, it is recommended to go through the official blog for more details.
@@ 3.1.6 Discussion
Combining different architectural components within SAM variants might further overcome the above deficiencies by complimentarily taking advantages from each architecture, leading to better balance between efficiency and accuracy. Intuitively, hybrid designs that integrate CNN-based structures with Transformer-based modules or Transformer-alternative models are feasible solutions. There have been attempts in recent SAM variants, such as Lite-SAM [fu2024lite], which employs a CNN-Transformer hybrid encoder LiteViT, and RWKV-SAM [yuan2024mamba], which integrates CNNs with the RWKV model. These variants with hybrid architectures have demonstrated significant improvements in efficiency while maintaining comparable segmentation performance to vanilla SAM. Despite their promise, the mismatch between heterogeneous layers, which may lead to low quality in feature representations and high difficulty in model optimization, is a non-negligible challenge for building up effective hybrid models, and as a result, more advanced designs about module combination are needed in the future research.
In addition to backbone selection, additional compression techniques such as quantization and pruning can further reduce model size and memory usage during inference. However, these methods may not always deliver the expected runtime speedup and can sometimes lead to a significant degradation in accuracy and generalization. The training procedure itself also has a profound impact on performance. Training from scratch can yield comprehensive, high-quality models but requires substantial training resources. Knowledge distillation offers a more efficient alternative using less data. Yet, distilled models may fail to fully inherit the teacher’s capabilities, resulting in performance inferior to models trained from scratch. Moreover, combinations of these methods may yield models with varied accuracy–efficiency profiles. For example, pruning combined with post-quantization can compress a model to an extremely lightweight variant by reducing both parameter size and model complexity; subsequent fine-tuning (either through full training or distillation) can then help recover lost accuracy, thereby achieving a more balanced design. The recent work DC-SAM [zheng2025deep] achieves deep compression on SAM by first leveraging distillation and subsequently applying quantization.
@@ 3.2 Accelerating SegEvery Tasks
As described in Section 3.1, the primary efficiency bottleneck for the SegAny task lies in the heavy image encoder. Any SAM variant with a more lightweight architecture will inherently be able to segment everything faster than the original SAM. However, as analyzed by Zhang et al. [zhang2023mobilesamv2], the main challenge of the SegEvery task stems from the dense grid sampling strategy. This strategy first predicts numerous masks based on a point grid and then selects valid masks, which is computationally expensive. Consequently, designing a more efficient sampling strategy to reduce the number of predicted masks has become the core approach to accelerating the SegEvery task. Another potential solution is to convert the SegEvery task into another well-established task, such as all-instance segmentation, as done in FastSAM [zhao2023fast]. In this part, we will review works that have specifically proposed optimized sampling strategies to accelerate the SegEvery task.
@@
Building on the structure of SAM, Zhang et al. [zhang2023mobilesamv2] introduced an object-aware prompt sampling strategy to enhance the efficiency of the SegEvery task. This project, named MobileSAMv2, operates independently of their previous work [zhang2023faster] focused on accelerating the SegAny task. In MobileSAMv2, the researchers employ the YOLOv8 model, trained on a small subset of SA-1B, for object discovery. This model generates a large number of bounding boxes corresponding to latent objects. Highly overlapping boxes are filtered using Non-Maximum Suppression (NMS), and the remaining boxes are used as box prompts. By using these filtered boxes as prompts, MobileSAMv2 eliminates the need to filter predicted masks—a much more time-consuming process. With the maximum number of prompts set to 320, the new strategy is reported to be 16 times faster than the traditional 32*32 grid sampling strategy. Additionally, MobileSAMv2 can be integrated with MobileSAM to create a unified model that achieves high efficiency in both SegAny and SegEvery tasks.
@@
Figure 14: The pipeline of hierarchical sampling strategy for TinySAM [shu2023tinysam]. The hierarchical strategy contains two rounds of sampling: one round for sparse sampling and the other round for dense sampling (including removing points from high confident regions). The final prediction comes from the fusion of results in two rounds.
@@
Meanwhile, Shu et al. [shu2023tinysam] also observed that using a dense points grid (e.g., 32*32, 64*64) often generates a large number of redundant masks, which are later filtered out during post-processing, an operation that incurs significant time costs. In reality, only a few points within the grid are necessary to produce confident masks. To address this inefficiency, they proposed a hierarchical strategy for efficient sampling, which progressively selects optimal points for mask generation. This strategy involves two rounds of prompt generation. In the first round, a sparse grid is used, incorporating only a fraction (approximately 1/4) of the default points along each side. Masks are generated based on these points, and after filtering, only high-confidence masks are retained as final predictions. In the second round, a denser grid is applied, following the default configuration. However, points located in regions already covered by confident masks are excluded, significantly reducing the number of points. The predicted results from both rounds are then fused to produce the final output. The pipeline for this hierarchical strategy is illustrated in Fig. 14. By employing this two-round approach, the sampling process becomes both more time-efficient and fine-grained, leading to a notable acceleration in the SegEvery task with minimal performance degradation.
@@
Figure 15: The architecture of Lite-SAM [fu2024lite]. It takes the Lite-ViT as an image encoder and proposes an automated prompt proposal network (AutoPPN) for efficient sampling, while the prompt encoder and the mask decoder are copied from the original SAM.
@@
Different from all aforementioned works, Fu et al. [fu2024lite] propose an end-to-end training pipeline specifically designed for the SegEvery task, aiming to develop a SAM variant capable of segmenting everything more efficiently. Their model, named Lite-SAM, retains the overall architecture of the original SAM but replaces the heavy image encoder with a more lightweight solution. The architecture overview of Lite-SAM is shown in Fig. 15. Lite-SAM employs a CNN-Transformer hybrid structure called Lite-ViT, which consists of four stages with 2, 2, 6, and 2 Lite-ViT blocks, respectively. The key innovation in Lite-ViT is the Multi-Scale Pooling Module (MSPM), which serves as an alternative to the traditional attention mechanism. Adapted from the PoolFormer block [yu2022metaformer], MSPM utilizes channel-wise LayerNorm and extends the pooling operation to multiple scales. As discussed earlier, another major bottleneck in SAM lies in the time-consuming grid sampling strategy. To address this, Lite-SAM introduces an Automated Prompt Proposal Network (AutoPPN) to improve sampling efficiency. AutoPPN takes the feature maps generated by the encoder as input and directly predicts point and box prompts. To ensure high-quality prompts, Lite-SAM uses a more powerful MSPM-based network instead of CNNs, and incorporates distance transform to estimate the confidence of point prompts. While Lite-SAM is primarily designed to accelerate the SegEvery task, it also demonstrates improved efficiency in the SegAny task due to its lightweight image encoder.
@@
Similar to Lite-SAM’s idea of automatic prompt generation, the recent variant AoP-SAM proposed by Chen et al. [chenaop], employs an efficient Prompt Predictor module to identify the optimal regions for prompt candidates, followed by an Adaptive Sampling and Filtering (ASF) module to generate prompts in a coarse-to-fine manner. The architecture of AoP-SAM is illustrated in Fig. 16. The Prompt Predictor consists of two encoder, processing original images and embedding from SAM’s image encoder respectively, and one decoder, incorporating the spatial and context information from encoders to generate the prompt confidence map. The function of ASF module is to extract essential prompts from prompt confidence map, which firstly samples the local maxima as coarse prompts, and secondly filters by the IoU scores of previously generated referenced masks and elimination scores of prompts, ensuring only crucial prompts will be retained for later mask generation. The elimination score is calculated by referenced mask and image embedding, indicating the possibility of being redundant. The experiments coudcted on three image segmentation datasets, have demonstrated AoP-SAM’s highly efficiency for SegEvery task, surpassing both SAM’s dense grid sampling strategy and MobileSAMv2’s [zhang2023mobilesamv2] object-aware sampling strategy.
@@
Figure 16: The architecture of AoP-SAM [chenaop]. It adopts a Prompt Predictor to predict prompt confidence map and an Adaptive Sampling and Filtering (ASF) module to generate final prompts.
@@ 3.3 Future Research Directions
Exploring Advanced Architectures. While current SAM variants have demonstrated efficiency gains through the adoption of efficient architectures and model compression techniques, there is substantial potential for further improvement. The exploration of Transformer-alternative models, such as Mamba [gu2023mamba], RetNet [sun2023retentive], KAN [liu2024kan], and TTT [sun2024learning], presents an exciting opportunity to design more lightweight and efficient structures. With the success of incorporating RWKV into SAM [yuan2024mamba], leveraging Transformer-alternative models as backbone has shown great potential to build up efficient SAM variants, which may offer advantages in computational efficiency without sacrificing segmentation accuracy. In addition to alternative models, refining attention modules’ efficiency remains critical for further development of accelerating SAM, inspired by the preliminary exploration of FastSAM3D [shen2024fastsam3d]. Methods such as linear attention, low-rank factorization, or hybrid architectures combining convolutional and attention-based designs, which have been widely applied to Transformer-based models [Han_2023_ICCV, hajimolahoseini2022strategies, fang2022hybrid], should be further investigated for efficient SAM. Addressing bottlenecks in both computation and memory usage will enhance SAM’s deployment across diverse hardware environments.
@@
Exploiting Sparsity and Acceleration Techniques. The inherent sparsity observed in deep neural networks, where only a subset of parameters significantly contributes to model output, offers a promising avenue for improving SAM’s efficiency. The discovery in very recent work, SuperSAM [abebe2025supersam], showing that the subnetwork with 7 layers in image encoder pruned (12 layers in total) can still maintain over 60% performance compared to vanilla SAM, further provides proofs to SAM’s sparsity, which poses a promising research direction for further uncovering the fine-grained sparsity in SAM’s architecture. Techniques like pruning, quantization, and structured sparsity [SAMfast] could further reduce SAM’s computational demands. While several initial efforts in sparsification, such as those leveraged in SlimSAM [chen20230] and PQ-SAM[liu2024pq], have shown success, future research could dive deeper into understanding the distribution and dynamics of sparsity in SAM. This includes investigating the optimal layers or components of SAM that can be pruned or sparsified without impacting performance [abebe2025supersam]. Additionally, techniques such as sparse attention mechanisms, dynamic pruning during inference, and low-precision training, which have been deployed on Transformer-based architectures [lou2024sparser, rao2021dynamicvit, NEURIPS2023_20bd42d8], can be further explored to balance accuracy and efficiency, especially for large-scale deployments. Recently, Fan et al. [fan2025compress] propose a novel compression technique, based on the Hyper-Compression principle, which projects high-dimensional parameters into scalars, providing new insights for lightening SAM. By combining these with advanced knowledge distillation techniques, more compact, efficient variants of SAM could be realized.
@@
Hardware-Specific Optimizations. Optimizing SAM for specific hardware platforms, including GPUs, TPUs, specialized AI accelerators (e.g., NVIDIA’s TensorRT or Google’s Edge TPU), and edge devices, can unlock significant gains in performance and efficiency. Equipping SAM with hardware-aware model optimization techniques, for instance, operator fusion, quantization-aware training, and custom CUDA kernels, some of which have been applied in SAMfast [SAMfast] by PyTorch extensions, requires further exploration in the future to maximize throughput and reduce latency when deploying SAM on modern hardware platforms. In the context of edge devices, which face extreme constraints in storage, computational power, and energy supply, these optimizations are crucial for real-time applications, such as segmentation on UAVs or IoT devices. Future research could explore hierarchical cloud-edge architectures to offload computationally expensive tasks to the cloud, while running lightweight models locally on edge devices. This idea has already been investigated in LLMs for robot control [luan2023hierarchical]. Additionally, leveraging specialized AI hardware like Field-Programmable Gate Arrays (FPGAs) [zeng2024flightllm] or using techniques such as hardware-aware neural architecture search (NAS) and mixed-precision quantization [xiao2023patch] might further optimize SAM for low-latency, resource-constrained environments, ensuring that the model operates effectively across diverse hardware platforms. Among them, NAS is now successfully applied to SAM, building up the first supernetwork of SAM [abebe2025supersam].
@@
Computation Cost Attacks and Defenses. Substantial progress has been made by the research community to improve the efficiency of deep learning models, particularly for deployment on resource-constrained edge devices. However, recent studies have exposed vulnerabilities in some models when subjected to computation cost attacks, which deliberately increase resource consumption and processing time far beyond normal operational levels [hong2020panda, chen2022nicgslowdown, chen2023dark, gao2024inducing, pan2022gradauto]. Despite these findings, the efficiency and robustness of Segment Anything Model (SAM) variants under computation cost attacks remain largely unexplored, especially when techniques like dynamic models are applied. This presents a valuable opportunity for future research. One direction could involve identifying specific vulnerabilities in SAM variants when exposed to such attacks. Additionally, the insights gained from these vulnerabilities could inform the development of defense strategies that ensure stable efficiency, particularly in resource-limited environments such as edge and mobile devices.
@@
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [ravi2024sam] for video segmentation and MM-SAM [xiao2024segment] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation [sener2020temporal] and cross-modal feature sharing [lu2020cross] could significantly reduce computational costs. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention [diba2023spatio], efficient memory mechanisms (which are now effective solutions for accelerating SAM 2 [xiong2024efficientsam2, zhou2025edgetam]) for temporal data, and early-fusion techniques [hemkerhybrid] to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.
@@
Efficient Variants of SAM 2. Segment Anything Model 2, successor of Segment Anything Model, which not only achieves higher accuracy and efficiency on image segmentation task but also extends its powerful capability into video segmentation tasks, has been increasingly gaining attention from industry and research community. However, similar to SAM’s dilemma, SAM 2 also suffers the inefficiency issue due to the large hierarchical image encoder, hindering its application in various resource-constrained, real-world scenarios. Moreover, another main bottleneck of SAM 2 lies in the newly introduced memory mechanism, more specifically, the memory attention module which attends feature map to memories, as memory features holds the same spatial size to image feature, resulting in long memory token sequence and heavy computational cost in cross-attention. Several works of efficient SAM 2 [liu2024surgical, xiong2024efficientsam2, zhou2025edgetam] has emerged focusing on solving these two bottlenecks. Surgical SAM 2 [liu2024surgical] has proposed the efficient frame pruning strategy which only retains the most informative frames in memory bank to achieve reduction in memory usage and acceleration in inference, providing a promising idea for efficient segmentation in videos. The subsequent variant, named EfficientTAM [xiong2024efficientsam2], has built up an efficient memory module, which reconstructs memories by firstly adopting average pooling in local window and secondly restoring the spatial size through repeating, leading to less computation in memory attention. Furthermore, EfficientTAM replaces the image encoder with SAMI [xiong2024efficientsam] pre-trained vanilla ViT, achieving higher efficiency and maintaining competing performance. Recently, a mobile-friendly SAM 2 variant, named EdgeTAM [zhou2025edgetam], further speeds up the memory attention by leveraging a global Perciever [jaegle2021perceiver] and a novel 2D Spatial Perceiver to effectively compress memories. Moreover, EdgeTAM opts for RepViT as backbone and retrains through a two-stage distillation pipeline with SAM 2 as teacher model, pursuing for efficiency and accuracy improvements. Quantization techniques have also been preliminarily employed in SAM 2, for instance, the newly proposed variant, Q-SAM2 [farronato2025q], which introduces calibration in SAM 2’s image encoder and subsequently implements quantization-aware training pipeline, effectively compressing model into low-bit configurations. Moreover, SAM 2 has also been rewritten by Pytorch Team [SAM2fast] with tricks like batching prompts and reducing precision, and techniques like AOTInductor’s (AOTI) ahead-of-time compilation, achieving noticeable enhancement on latency. Compared to SAM, there are much less works on efficient variants for SAM 2 up till now, posing it a challenging task. Techniques and strategies proved to be effective in accelerating SAM, have not been fully investigated on SAM 2, which might be the potential direction for future research. For the newly published SAM 3 [carion2025sam3segmentconcepts], which generalized from its predecessor by combining a DETR-based detector with the SAM 2 style tracker, suffers similar efficiency bottlenecks as SAM 2 does. And the recently proposed method, EfficientSAM3 [zeng2025efficientsam3], has targeted at SAM3’s inefficiency in its shared vision encoder and dense memory tracker. Two stages of distillation are leveraged to progressively lighten SAM 3, with a final stage of end-to-end fine-tuning for joint model refinement.